OpenStudy (anonymous):

methyl salicylate has mass percent composition of 63.2% C, 31.6% O and 5.26% H. The mass of one molecule of the compound is 2.53 x 10^-22 g. Determine the empirical formula and the molecular formula of methyl salicylate.

OpenStudy (anonymous):

to find the EF first find the ratio of n of moles in each compound: since the % is given we know that in 100g of methyl salicylate, we will have: 63.2g of C , 31.6g of O and 5.26g of H n=m/M so n(c)=63.2/12.01 n(O)=31.6/16 n(H)=5.26/1.008 =5.26 mol =1.975 mol =5.218 mol ratio for C : O : H >> 2.66 : 1.975 : 2.642 divide everything by the smallest number to get the simplest ratio, we get >> 2.66 : 1 : 2.642 times everything by 3 to get whole number >> 8 : 3 : 8 (round off) so EF= C8H8O3 m(EF)= 8x12.01+8x1.008+3x16 =152.144g m(MF)= 6.022x10^23 x 2.53x10^-22 (avagadro's number (number of atoms in one mole) times the mass of one atom) =153.36g thus MF=EF MF=C8H8O3